Ross Co., Westerfield, Inc., and Jordan Company announced a new agreement to market their respective products in China on July 18 (7/18), February 12 (2/12), and October 7 (10/7), respectively. Given the information below, calculate the cumulative abnormal return (CAR) for these stocks as a group. Assume all companies have an expected return equal to the market return. (A negative value should be indicated by a minus sign. Leave no cells blank - be certain to enter "0" wherever required. Do not round intermediate calculations. Round your answers to 1 decimal place.)

Ross Co. Westerfield, Inc. Jordan Company
Date Market
Return
Company
Return
Date Market
Return
Company
Return
Date Market
Return
Company
Return
7/12 -0.4 -0.5 2/8 -0.4 -0.5 10/1 2.1 0.9
7/13 1.9 0.3 2/9 -0.5 -0.5 10/2 2 0.9
7/16 -0.3 0.2 2/10 0.2 0.6 10/3 0.5 1.3
7/17 -0.6 -0.2 2/11 0.4 2.8 10/6 -0.1 -0.1
7/18 -0.1 1.3 2/12 -0.1 0.1 10/7 -1.5 -0.1
7/19 -2.6 0 2/15 1.2 2 10/8 1.2 0.5
7/20 -0.9 -0.1 2/16 0.7 0.7 10/9 -0.5 -0.4
7/23 0.6 0.4 2/17 -0.1 0 10/10 0.1 -0.1
7/24 1.9 0 2/18 2.3 0.4 10/13 -0.4 -0.6
Answer #1

Abnormal return is calculated as compnay return in excess of expected return. Since all companies have an expected return equal to the market return, hence

Abnormal return = Company return - market return.

We will use this formula to calculate the abnormal return for each of the three compnaies. We will then add to get the sum, divide the sum to get the average and then cumulate the average to get CAR.

Please see the table below. The last column is your answer.

Ross Co. Westerfield, Inc. Jordan Company Abnormal Returns Sum Average CAR
Date Market Company Date Market Company Date Market Company Days from Ross Co. Westerfield Jordan
Return Return Return Return Return Return announcement
A B C D E F G = B - A H = D - C I = F - E J = G + H + I K = J / 3 Cumulative of K
12-Jul -0.4 -0.5 8-Feb -0.4 -0.5 1-Oct 2.1 0.9 -4 -0.1 -0.1 -1.2 -1.4 -0.5 -0.5
13-Jul 1.9 0.3 9-Feb -0.5 -0.5 2-Oct 2 0.9 -3 -1.6 0 -1.1 -2.7 -0.9 -1.4
16-Jul -0.3 0.2 10-Feb 0.2 0.6 3-Oct 0.5 1.3 -2 0.5 0.4 0.8 1.7 0.6 -0.8
17-Jul -0.6 -0.2 11-Feb 0.4 2.8 6-Oct -0.1 -0.1 -1 0.4 2.4 0 2.8 0.9 0.1
18-Jul -0.1 1.3 12-Feb -0.1 0.1 7-Oct -1.5 -0.1 0 1.4 0.2 1.4 3 1.0 1.1
19-Jul -2.6 0 15-Feb 1.2 2 8-Oct 1.2 0.5 1 2.6 0.8 -0.7 2.7 0.9 2.0
20-Jul -0.9 -0.1 16-Feb 0.7 0.7 9-Oct -0.5 -0.4 2 0.8 0 0.1 0.9 0.3 2.3
23-Jul 0.6 0.4 17-Feb -0.1 0 10-Oct 0.1 -0.1 3 -0.2 0.1 -0.2 -0.3 -0.1 2.2
24-Jul 1.9 0 18-Feb 2.3 0.4 13-Oct -0.4 -0.6 4 -1.9 -1.9 -0.2 -4 -1.3 0.9